Direct Labor Variances

Bellingham Company produces a product that requires 10 standard hours per unit at a standard hourly rate of $20.00 per hour. If 2,900 units required 29,600 hours at an hourly rate of $19.00 per hour, what is the direct labor (a) rate variance, (b) time variance, and (c) total direct labor cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Direct labor rate variance $ U or F
b. Direct labor time variance $ U or F
c. Total direct labor cost variance $ U or F
Answer #1

Labor rate variance = (SR - AR) * AH

= (20 - 19)*29600

= -29,600 F (minus sign)

Labor time variance = (SH - AH) * SR

= (2900*10 - 29600) * 20

= 12,000 U

Labor cost variance = 29600 F + 12000 U

= -17,600 F